OpenStudy (anonymous):

$4^{\frac{1}{x}-2}= \frac{(ln \sqrt{e})}{2}$

OpenStudy (whpalmer4):

\[4^{\frac{1}{x}-2}= \frac{(ln \sqrt{e})}{2}\] (If you put \ [ and \ ] (no spaces) instead of $ around your expressions it should format them)

OpenStudy (whpalmer4):

Why not take the log of both sides, then solve for x?

OpenStudy (whpalmer4):

Remember that \[\log a^n = n \log a\]

OpenStudy (whpalmer4):

Also, try writing that radical sign with exponential notation instead and see if any ideas come to you...

OpenStudy (whpalmer4):

Don't forget that \(a^{-n} = \dfrac{1}{a^n}\)